OpenStudy (anonymous):

Evaluate the expression for the given of the variable(s). -(4h-1)/1+h ; h = 2

OpenStudy (ivettef365):

so you have -(4h-1) ------- 1 + h

OpenStudy (ivettef365):

just substitute the h for 2 as follows: -(4(2)-1) -------- 1 + 2 now you can solve

OpenStudy (anonymous):

So its 9?

OpenStudy (ivettef365):

you have -(8-1) = -(7) = -7/3

OpenStudy (anonymous):

Now im confused. How is it -7/3

OpenStudy (ivettef365):

you have a division as follows: -(4(2)-1) -(8-1) -(7) -7 -------- = -------- = ----- = ---- 2+1 3 3 3

OpenStudy (anonymous):

Oh okay, Thank you so much.

OpenStudy (ivettef365):

yw :)